OpenStudy (anonymous):

Solve x2 + 4x + 8 = 0

OpenStudy (anonymous):

\[x^2+4x+8=0?\]

OpenStudy (anonymous):

use quadratic formula solve\[\frac{-b \pm \sqrt{b^2-(4*a*c)}}{2a}\] a= 1 b=4 c=8

OpenStudy (anonymous):

x=-2+2i or x=-2-2i